# On the slow phase for fixed-energy Activated Random Walks

### v1

@@ 5 Upper bound for the number of frozen particles
The main goal of this Section is to prove Proposition 4.3. In this section, the analysis is restricted to the dynamics of a single mode.
@@ 5.1 $\sigma$-algebras and flow of particles
To each site $y\in[iK,iK+a]$, $i=0,1,\dots,n+1$ we associate a i.i.d. stack of instructions, $\xi^{y}=(\xi^{y}_{k})_{k\in\mathbb{N}}$, as explained in Section 2. Each time we topple one of those sites we examine the next unused instruction in $\xi^{y}$. For sites $y\in[iK+a+1,(i+1)K)$ we associate two independent copies of the stack of instructions, $\xi^{y,L}$ and $\xi^{y,R}$. Each time one of those sites is toppled we use the first unused instruction in $\xi^{y,L}$ if the hot particle is chosen at $\{iK,iK+a\}$ or the first unused instruction in $\xi^{y,R}$ if the hot particle is chosen at $\{(i+1)K,(i+1)K+a\}$. That is, we use $\xi^{y,R}$ for particles coming “from the right” and $\xi^{y,L}$ for particles coming “from the left”.
@@
$\mathcal{F}_{i}=\sigma(\{\xi^{y}:y\in(0,iK+a]\}\cup\{\xi^{x,L}:x\in(iK+a,(i+1)K)\}).$
Thus, $\mathcal{F}_{i}$ comprises information from particles being emitted from block $i$ to block $i+1$ but no information from particles going of block $i+1$ to block $i$.
Let $\eta$ be the configuration of the ring at the beginning of a mode, and, for $0\leq k\leq n$, define $\eta_{k}$ to be the restriction of $\eta$ to blocks $0,1,\dots,k$. Let $d^{\prime}$ be the number of defects in $\eta_{k}$. We define, for $(m,d)\in\{0\}\times\{0,1,\dots,d^{\prime}\}\cup\mathbb{N}\times\{d^{\prime}\}$, $\eta^{\prime}_{k}(m,d)$ in the following way: first, we use the toppling procedure of Section 3 restricted to blocks $0,1,\dots k$. After this is done, we define $\eta^{\prime}_{k}(m,d)$ as the resulting configuration after block $k$ receives $m+d$ particles from block $k+1$ ($d$ particles for defective sites, $m-1$ particles arrive necessarily at $kK+a$ and $1$ particle may arrive either at $kK$ if $kK$ is vacant and there are defects in $(kK-K/2,kK)$ or at $kK+a$ if not.)
Let $1\leq i\leq k\leq n$, and $m$ and $d$ integers for which $\eta^{\prime}_{k}(m,d)$ is well-defined. We run the procedure described on Section 3 on $\eta^{\prime}_{k}(m,d)$ and define the following random functions:
  • $L_{i}^{k}(m,d)$ is the number of particles that are emitted from block $i$ to block $i-1$;
@@
By (P7), $S_{i}^{k}(m,d)$ can only be 00 or $1$.
We also define $M^{k}_{i}$ as the number of particles emitted from block $i$ to block $i-1$ that will be available for attempted emissions in block $i-1$.
@@
A particle that is emitted $i$ to block $i-1$ can only fix a defect in block $i-1$, reach the vacant hole or wait at $iK+a$ to be used in an attempted emission. Therefore, we can state the following mass-balance equation
$L_{i}^{k}(m_{i},d_{i})=M_{i-1}^{k}+D_{i-1}^{k}.$ (5.1)
@@
For $i=n$, the result is immediate. Fix some $i\in\{1,\dots,n-1\}$. Whenever possible, the procedure described on Section 3 chooses the hot particle in blocks $11,2,\dots,i-$ 1, and when this choice is possible, it depends only on the configuration in blocks $11,2,\dots,i-$ 1. When the choice in those blocks is not possible, the procedure either stops or waits for a particle to come to block $i$ from block $i+1$. This particle will either fix a defect in block $i$, reach the vacant hole at $iK$ or wait at $iK+a$ until it is chosen to be the hot particle. By the way the procedure was designed, these particles will have the least priority to be chosen as the hot particle, and will not affect the dynamics in the first $i$ blocks until they are chosen to be hot.
@@ 5.2 Bounding the exponential expectation
To prove Proposition 4.3, we will use the following one-block estimate, to be proved in Section 6.
Proposition 5.2.
@@
    For any initial configuration $\eta_{0}$ satisfying Properties (P1) to (P9),
    $\sup_{\ell\geq 0,d\geq 0}\mathbb{E}\Bigg{[}\sum_{m=0}^{\infty}e^{48S_{i}(m,d)}\mathds{1}_{\{L_{i}(m,d)=\ell\}}\mathds{1}_{\{L_{i}(m,d)+R_{i}(m,d)\geq 1\}}~{}\Bigl{|}~{}\mathcal{F}_{i-1}\Bigg{]}\leq e^{4}.$
@@
    For any initial configuration $\eta_{0}$ satisfying Properties (P1) to (P9) and with holes only at positions $iK$, $1\leq i\leq n$,
    $\sup_{\ell\geq 0,d\geq 0}\mathbb{E}\Bigg{[}\sum_{m=0}^{\infty}e^{48S_{i}(m,d)}\mathds{1}_{\{L_{i}(m,d)=\ell\}}~{}\Bigl{|}~{}\mathcal{F}_{i-1}\Bigg{]}\leq e^{4},$
@@
By definition, $F(E_{n})$ is the number of particles that are frozen on the collection of blocks that perform at least one emission. Also, at the beginning of the procedure there are at most $n$ free particles (one at each block), and the number of free particles never increases.Since $\zeta>1-1/4K$, we can bound the number of defects by $n$. Therefore, using the definition of $S_{i}$, Proposition 5.1 and the mass-balance equation (5.1):
$\mathbb{E}[e^{48F(E_{n})}]$ $=\mathbb{E}\Bigg{[}\sum_{\begin{subarray}{c}d_{0},\dots d_{n}\\ m_{0},\dots,m_{n}\end{subarray}}\prod_{i=1}^{n}e^{48S_{i}(m_{i},d_{i})}\mathds{1}_{\{M_{i}^{n}=m_{i}\}}\mathds{1}_{\{D_{i}^{n}=d_{i}\}}\mathds{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}+d_{i-1}\}}\mathds{1}_{G^{i}_{m_{i},d_{i}}}\Bigg{]}$
$\leq\sum_{\begin{subarray}{c}d_{0},\dots d_{n}\\ d_{0}+\dots+d_{n}\leq n\end{subarray}}\mathbb{E}\Bigg{[}\sum_{m_{0}=0}^{n}\sum_{m_{1},\dots,m_{n}}\prod_{i=1}^{n}e^{48S_{i}(m_{i})}\mathds{1}_{\{L_{i}(m_{i})=m_{i-1}+d_{i-1}\}}\mathds{1}_{G^{i}_{m_{i},d_{i}}}\Bigg{]}.$
@@
$\mathbb{E}[e^{48F(E_{n})}]\leq 4^{n}\cdot\sup_{\{d_{0},\dots,d_{n}\}}\mathbb{E}\Bigg{[}\sum_{m_{0}=0}^{n}\sum_{m_{1},\dots,m_{n-1}}\prod_{i=1}^{n}e^{48S_{i}(m_{i})}\mathds{1}_{\{L_{i}(m_{i})=m_{i-1}+d_{i-1}\}}\mathds{1}_{G^{i}_{m_{i},d_{i}}}\Bigg{]}.$
@@
[…]
Indeed, for $r=1$ the expectation above is bounded by $n+1$, and the base case of induction is finished. Suppose the inequality is valid for every integer between $1$ and $r-1$. By conditioning on $\mathcal{F}_{r-1}$ and using item (i) of Proposition 5.2,
$\!\!\!\!\!\!\!\!\!\!\!\!\mathbb{E}\Bigg{[}\sum_{m_{0}=0}^{n}\cdots\sum_{m_{r}}\prod_{i=1}^{r}e^{48S_{i}(m_{i},d_{i})}\mathds{1}_{\{L_{i}(m_{i})=m_{i-1}+d_{i-1}\}}\mathds{1}_{G^{i}_{m_{i},d_{i}}}\Bigg{]}$
$\leq\mathbb{E}\Bigg{[}\sum_{m_{0}}\sum_{m_{1}}\cdots\sum_{m_{r-1}}\prod_{i=1}^{r-1}e^{48S_{i}(m_{i},d_{i})}\mathds{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}+d_{i-1}\}}\mathds{1}_{G^{i}_{m_{i},d_{i}}}\Bigg{]}$
$\qquad\cdot\sup_{\ell\geq 0,d\geq 0}\mathbb{E}\Bigg{[}\sum_{m_{r}}e^{48S_{r}(m_{r},d)}\mathds{1}_{\{L_{r}(m_{r},d)=\ell\}}\mathds{1}_{G^{r}_{m_{r},d_{r}}}\mid\mathcal{F}_{r-1}\Bigg{]}$
$\leq e^{4}\mathbb{E}\Bigg{[}\sum_{m_{0}}^{n}\cdots\sum_{m_{r-1}}\prod_{i=1}^{r-1}e^{48S_{i}(m_{i},d_{i})}\mathds{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}\}}\mathds{1}_{G^{i}_{m_{i},d_{i}}}\Bigg{]}$
@@
Proof of Proposition 4.4.
The proof of this proposition is completely analogous to the proof of Proposition 4.3, but summing over all the blocks, not only the ones that realize at least one emission, and using item $(ii)$ of Proposition 5.2, because in the first mode all holes are at positions $iK$, $i=1,2,.\dots,n$. ∎
@@ 6 One-block estimate
This section is devoted to proving Proposition 5.2. We will begin by proving item $(i)$.
Let $i$ be fixed. We will consider two cases. First, we will analyze the case where there are defects in $[iK-K/2,iK+K/2)$ after we stabilize $\eta^{\prime}_{i}(m,d)$ using the procedure described in Section 3. The only value of $m$ compatible to the definition of $\eta^{\prime}_{k}$ in this case is $m=0$, and, by Properties (P3) and (P8), $S_{i}(0,d)=0$. It follows that
$\sup_{\ell\geq 0,d\geq 0}$ $\mathbb{E}\Bigg{[}\sum_{m=0}^{\infty}e^{48S_{i}(m,d)}\mathds{1}_{\{L_{i}(m,d)=\ell\}}\mathds{1}_{G^{i}_{m,d}}\mid\mathcal{F}_{i-1}\Bigg{]}$
$=\sup_{\ell\geq 0,d\geq 0}\mathbb{E}[e^{48S_{i}(0,d)}\mathds{1}_{\{L_{i}(0,d)=\ell\}}\mathds{1}_{G^{i}_{0,d}}\mid\mathcal{F}_{i-1}]\leq 1.$
We now analyze the remaining case, i.e., when there are no defects in $[iK-K/2,(i+1)K/2)$ after we stabilize $\eta^{\prime}_{i}(m,d)$ using the procedure described in Section 3. Thus, we need only to analyze the case where $d=0$.
We will use from now on a slight abuse of notation and denote by $L_{i}(j),S_{i}(j)$ and $\mathds{1}_{G^{i}_{j}}$ the values of the quantities $L_{i},S_{i}$ and $\mathds{1}_{G^{i}}$ after $j$ attempted emissions in block $i$.
For each particle added at $iK+a$, at least one attempted emission will be made. In other words, for each $m$ there is at least one corresponding $j$. Therefore,
$\sum_{m=0}^{\infty}e^{48S_{i}(m,d)}\mathds{1}_{\{L_{i}(m,d)=\ell\}}\mathds{1}_{G^{i}_{m,d}}\leq\sum_{j=0}^{\infty}e^{48S_{i}(j)}\mathds{1}_{\{L_{i}(j)=\ell\}}\mathds{1}_{G^{i}_{j}}.$
@@
$\sup_{\ell\geq 0}\mathbb{E}\Bigg{[}\sum_{j=0}^{\infty}e^{48S_{i}(j)}\mathds{1}_{\{L_{i}(j)=\ell\}}\mathds{1}_{G^{i}_{j}}\mid\mathcal{F}_{i-1}\Bigg{]}\leq e^{4}.$ (6.1)
@@
It is immediate that $H(j)$ is $\mathcal{G}_{j}$-measurable. For brevity we will define ~$\tilde{\mathbb{P}}[\cdot]:=\mathbb{P}[\cdot\mid\mathcal{F}_{i-1}]$ and $\tilde{\mathbb{E}}[\cdot]:=\mathbb{E}[\cdot\mid\mathcal{F}_{i-1}]$. To prove Proposition 5.2 we will make use of the following lemmas.
Lemma 6.2.
@@
~$\tilde{\mathbb{P}}[H(j)>a/2\mid\mathcal{G}_{j-1}]\mathds{1}_{\{H(j-1)\in[0,a/2]\cup\{a\}\}}\leq e^{-150},$ (6.2)
~$\tilde{\mathbb{P}}(a/2<H(j)<a\mid\mathcal{G}_{j-1})<e^{-150}.$ (6.3)
Remark 6.3.
@@
Lemma 6.4.
@@
~$\tilde{\mathbb{P}}(L(j+2)>L(j)\mid\mathcal{G}_{j})\geq\frac{1}{5}.$
For every integer $\ell\geq 0$, denote by $\tau_{\ell}$ the first $j$ for which $L(j)=\ell$, with $\tau_{-1}:=-1$. We have that $\tau_{\ell+1}>\tau_{\ell}$, because every attempted emission will result in at most one particle emitted to the left. We can rewrite (6.1) to obtain
$\mathbb{E}$ $\left[\sum_{j=\tau_{\ell-1}+1}^{\infty}e^{48S(j)}\mathds{1}_{\{L(j)=\ell\}}\mathds{1}_{G_{j}}\mid\mathcal{F}_{i-1}\right]$
$\leq\sum_{j=\tau_{\ell-1}+1}^{\infty}\tilde{\mathbb{E}}[\mathds{1}_{\{L(j)=\ell\}}\mathds{1}_{\{S(j)=0\}}+e^{48}\mathds{1}_{\{L(j)=\ell\}}\mathds{1}_{G_{j}}\mathds{1}_{\{S(j)=1\}}]$
$\leq\tilde{\mathbb{E}}[\tau_{\ell+1}-\tau_{\ell}]+\sum_{k=1}^{\infty}\tilde{\mathbb{E}}[e^{48}\mathds{1}_{\{L(\tau_{\ell-1}+k)=\ell\}}\mathds{1}_{G_{\tau_{\ell-1}+k}}\mathds{1}_{\{S(\tau_{\ell-1}+k)=1\}}].$ (6.4)
Using Lemma 6.4, we get
$\tilde{\mathbb{E}}[\tau_{\ell+1}-\tau_{\ell}]=\sum_{r=0}^{\infty}\tilde{\mathbb{P}}(\tau_{\ell+1}-\tau_{\ell}>r)\leq\sum_{r=0}^{\infty}\left(\frac{4}{5}\right)^{\lfloor r/2\rfloor}<10.$ (6.5)
We split the sum in (6.4) at $k_{0}:=250$. By Lemma 6.4,
$\sum_{k=k_{0}}^{\infty}$ $\tilde{\mathbb{E}}[e^{48}\mathds{1}_{\{L(\tau_{\ell-1}+k)=\ell\}}\mathds{1}_{G_{\tau_{\ell-1}+k}}\mathds{1}_{\{S(\tau_{\ell-1}+k)=1\}}]$
$\leq\sum_{k=k_{0}}^{\infty}e^{48}\tilde{\mathbb{P}}(L(\tau_{\ell-1}+k)=\ell)\leq\sum_{k=k_{0}}^{+\infty}e^{48}\left(\frac{4}{5}\right)^{\lfloor k/2\rfloor}<1.$ (6.6)
@@
$\tilde{\mathbb{E}}[\mathds{1}_{\{L(\tau_{\ell-1}+k)=\ell\}}\mathds{1}_{G_{\tau_{\ell-1}+k}}\mathds{1}_{\{S(\tau_{\ell-1}+k)=1\}}]<k_{0}^{-1}e^{-48}.$ (6.7)
Let us analyze first the case $\ell\geq 1$. In this case, obviously $\mathds{1}_{G_{\tau_{\ell-1}}}=1$. First, we consider $k=1$. Note that if $S(\tau_{\ell-1}+1)=1$ and $L(\tau_{\ell-1}+1)=\ell$, then the $(\tau_{\ell-1}+1)$-th attempted emission is successful and $H(\tau_{\ell-1})=a$. Using (6.2), after conditioning on $\mathcal{G}_{\ell-1}$,
$\tilde{\mathbb{E}}$ $[\mathds{1}_{\{L(\tau_{\ell-1}+1)=\ell\}}\mathds{1}_{\{S(\tau_{\ell-1}+1)=1\}}]$
$\leq\tilde{\mathbb{E}}[\mathds{1}_{\{L(\tau_{\ell-1}+1)=\ell\}}\mathds{1}_{\{H(\tau_{\ell-1}+1)=a\}}\mathds{1}_{\{H(\tau_{\ell-1})=a\}}]$

### v2

@@ 6 Upper bound for the number of frozen particles
The main goal of this Section is to prove Proposition 3. In this section, the analysis is restricted to the dynamics of a single mode.
@@ 6.1 $\sigma$-algebras and flow of particles
To each site $y\in[iK-K/2,iK+K/2)$, $i=0,i,\dots,n+1$ we associate three independent copies of the stack of instructions, $\xi^{y}$, $\xi^{y,L}$ and $\xi^{y,R}$. Each time one of those sites is toppled we use the first unused instruction in $\xi^{y,L}$ if the hot particle is chosen at $\{(i-1)K,(i-1)K+a\}$, the first unused instruction in $\xi^{y}$ if the hot particle is chosen at $\{iK,iK+a\}$ or the first unused instruction in $\xi^{y,R}$ if the hot particle is chosen at $\{(i+1)K,(i+1)K+a\}$. Thus, each site $y$ in block $i$ has a stack associated to the block i1i-1 (except for $i=0$), one stack associated to block $i$ and one stack associated to block i1i-1 (except for $i=n+1$).
@@
$\mathcal{F}_{i}=\sigma$ $(\{\xi^{y},\xi^{y,R},\xi^{y,L}:y\in(-K/2,iK-K/2)\}$
$\cup\{\xi^{y,L},\xi^{y}:y\in[iK-K/2,iK+K/2)\}$
$\cup\{\xi^{y,L}:y\in[iK+K/2,(i+1)K+K/2)\}).$
In words, $\mathcal{F}_{i}$ comprises information about stacks associated with blocks $0,1,\dots,i$.
Let $\eta$ be the configuration of the ring at the beginning of a mode, and, for $0\leq k\leq n$, define $\eta_{k}$ to be the restriction of $\eta$ to all the blocks $0,1,\dots,k$. Let $d^{\prime}$ be the number of defects in $\eta_{k}$. We define, for $(m,d)\in\{0\}\times\{0,1,\dots,d^{\prime}\}\cup\mathbb{N}\times\{d^{\prime}\}$, $\eta^{\prime}_{k}(m,d)$ in the following way: first, we use the toppling procedure of Section 3 restricted to blocks $0,1,\dots k$. After this is done, we define $\eta^{\prime}_{k}(m,d)$ as the resulting configuration after block $k$ receives $m+d$ particles from block $k+1$ ($d$ particles for defective sites, m1m-1 particles arrive necessarily at $kK+a$ and $1$ particle may arrive either at $kK$ if $kK$ is vacant and there are defects in (kKK/2,kK)(kK-K/2,kK) or at $kK+a$ if not.)
Let $1\leq i\leq k\leq n$, and $m$ and $d$ integers for which $\eta^{\prime}_{k}(m,d)$ is well-defined. We run the procedure described in Section 3 on $\eta^{\prime}_{k}(m,d)$ and define the following random functions:
  • $L_{i}^{k}(m,d)$ is the number of particles that are emitted from block $i$ to block i1i-1;
@@
By (P7), $S_{i}^{k}(m,d)$ can only be 0 or $1$.
We also define $M^{k}_{i}$ as the number of particles emitted from block $i$ to block i1i-1 that will be available for attempted emissions in block i1i-1.
@@
A particle that is emitted from block $i$ to block i1i-1 can only fix a defect in block i1i-1, reach the vacant hole or wait at $iK+a$ to be used in an attempted emission. Therefore, we can state the following mass-balance equation
$L_{i}^{k}(M_{i}^{k},D_{i}^{k})=M_{i-1}^{k}+D_{i-1}^{k}.$ (1)
@@
For $i=n$, the result is immediate. Fix some $i\in\{1,\dots,n-1\}$. Whenever possible, the procedure described on Section 3 chooses the hot particle in blocks $1,2,\dots,i-1$, and when this choice is possible, it depends only on the configuration in blocks $1,2,\dots,i-1$. When the choice in those blocks is not possible, the procedure either stops or waits for a particle to come to block $i$ from block $i+1$. This particle will either fix a defect in block $i$, reach the vacant hole at $iK$ or wait at $iK+a$ until it is chosen to be the hot particle. By the way the procedure was designed, these particles will have the least priority to be chosen as the hot particle, and will not affect the dynamics in the first $i$ blocks until they are chosen to be hot.
@@ 6.2 Bounding the exponential expectation
To prove Proposition 3, we will use the following one-block estimate, to be proved in Section 7.
Proposition 6.
@@
    For any initial configuration $\eta_{0}$ satisfying Properties (P1) to (P9) and $K$ large enough,
    $\sup_{\ell\geq 0,d\geq 0}\mathbb{E}\Bigg[\sum_{m=0}^{\infty}e^{48S_{i}(m,d)}\mathbf{1}_{\{L_{i}(m,d)=\ell\}}\mathbf{1}_{\{L_{i}(m,d)+R_{i}(m,d)\geq 1\}}~\Bigl|~\mathcal{F}_{i-1}\Bigg]\leq e^{4}.$
@@
    For any initial configuration $\eta_{0}$ satisfying Properties (P1) to (P9) and with holes only at positions $iK$, $1\leq i\leq n$ and $K$ large enough,
    $\sup_{\ell\geq 0,d\geq 0}\mathbb{E}\Bigg[\sum_{m=0}^{\infty}e^{48S_{i}(m,d)}\mathbf{1}_{\{L_{i}(m,d)=\ell\}}~\Bigl|~\mathcal{F}_{i-1}\Bigg]\leq e^{4},$
@@
By definition, $F(E_{n})$ is the number of particles that are frozen on the collection of blocks that perform at least one emission. Also, at the beginning of the procedure there are at most $n$ free particles (one at each block), and the number of free particles never increases. Since $\zeta>1-1/(4K)$, we can bound the number of defects by $n$. Therefore, using the definition of $S_{i}$, Proposition 5 and the mass-balance equation (1):
$\mathbb{E}[e^{48F(E_{n})}]$ $=\mathbb{E}\Bigg[\sum_{\begin{subarray}{c}d_{0},\dots d_{n}\\ m_{0},\dots,m_{n}\end{subarray}}\prod_{i=1}^{n}e^{48S_{i}(m_{i},d_{i})}\mathbf{1}_{\{M_{i}^{n}=m_{i}\}}\mathbf{1}_{\{D_{i}^{n}=d_{i}\}}\mathbf{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}+d_{i-1}\}}\mathbf{1}_{G^{i}_{m_{i},d_{i}}}\Bigg]$
$\leq\sum_{\begin{subarray}{c}d_{0},\dots d_{n}\\ d_{0}+\dots+d_{n}\leq n\end{subarray}}\mathbb{E}\Bigg[\sum_{m_{0}=0}^{n}\sum_{m_{1},\dots,m_{n}}\prod_{i=1}^{n}e^{48S_{i}(m_{i},d_{i})}\mathbf{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}+d_{i-1}\}}\mathbf{1}_{G^{i}_{m_{i},d_{i}}}\Bigg].$
@@
$\mathbb{E}[e^{48F(E_{n})}]\leq 4^{n}\cdot\sup_{\{d_{0},\dots,d_{n}\}}\mathbb{E}\Bigg[\sum_{m_{0}=0}^{n}\sum_{m_{1},\dots,m_{n}}\prod_{i=1}^{n}e^{48S_{i}(m_{i},d_{i})}\mathbf{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}+d_{i-1}\}}\mathbf{1}_{G^{i}_{m_{i},d_{i}}}\Bigg].$
@@
$\mathbb{E}\Bigg[\sum_{m_{0}=0}^{n}\sum_{m_{1},\dots,m_{r}}\prod_{i=1}^{r}e^{48S_{i}(m_{i},d_{i})}\ \mathbf{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}+d_{i-1}\}}\mathbf{1}_{G^{i}_{m_{i},d_{i}}}\Bigg]\leq(n+1)e^{4r}.$
Indeed, for $r=1$ the expectation above is bounded by $n+1$, and the base case of induction is finished. Suppose the inequality is valid for every integer between $1$ and r1r-1. By conditioning on $\mathcal{F}_{r-1}$ and using item (i) of Proposition 6,
$\!\!\!\!\!\!\!\!\!\!\!\!\mathbb{E}\Bigg[\sum_{m_{0}=0}^{n}\cdots\sum_{m_{r}}\prod_{i=1}^{r}e^{48S_{i}(m_{i},d_{i})}\mathbf{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}+d_{i-1}\}}\mathbf{1}_{G^{i}_{m_{i},d_{i}}}\Bigg]$
$\leq\mathbb{E}\Bigg[\sum_{m_{0}}\sum_{m_{1}}\cdots\sum_{m_{r-1}}\prod_{i=1}^{r-1}e^{48S_{i}(m_{i},d_{i})}\mathbf{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}+d_{i-1}\}}\mathbf{1}_{G^{i}_{m_{i},d_{i}}}\Bigg]$
$\hskip 17.00024pt\cdot\sup_{\ell\geq 0,d\geq 0}\mathbb{E}\Bigg[\sum_{m_{r}}e^{48S_{r}(m_{r},d)}\mathbf{1}_{\{L_{r}(m_{r},d)=\ell\}}\mathbf{1}_{G^{r}_{m_{r},d_{r}}}\mid\mathcal{F}_{r-1}\Bigg]$
$\leq e^{4}\mathbb{E}\Bigg[\sum_{m_{0}}^{n}\cdots\sum_{m_{r-1}}\prod_{i=1}^{r-1}e^{48S_{i}(m_{i},d_{i})}\mathbf{1}_{\{L_{i}(m_{i},d_{i})=m_{i-1}\}}\mathbf{1}_{G^{i}_{m_{i},d_{i}}}\Bigg]$
@@
Proof of Proposition 4.
The proof of this proposition is completely analogous to the proof of Proposition 3, but summing over all the blocks, not only the ones that realize at least one emission, and using item $(ii)$ of Proposition 6, because in the first mode all holes are at positions $iK$, $i=1,2,.\dots,n$. ∎
@@ 7 One-block estimate
This section is devoted to proving Proposition 6. We will begin by proving item $(i)$.
Let $i$ be fixed. We will consider two cases. First, we will analyze the case where there are defects in [iKK/2,iK+K/2)[iK-K/2,iK+K/2) after we stabilize $\eta^{\prime}_{i}(m,d)$ using the procedure described in Section 3. The only value of $m$ compatible with the definition of $\eta^{\prime}_{k}$ in this case is $m=0$, and, by Properties (P3) and (P8), $S_{i}(0,d)=0$. It follows that
$\sup_{\ell\geq 0,d\geq 0}$ $\mathbb{E}\Bigg[\sum_{m=0}^{\infty}e^{48S_{i}(m,d)}\mathbf{1}_{\{L_{i}(m,d)=\ell\}}\mathbf{1}_{G^{i}_{m,d}}\mid\mathcal{F}_{i-1}\Bigg]$
$=\sup_{\ell\geq 0,d\geq 0}\mathbb{E}[e^{48S_{i}(0,d)}\mathbf{1}_{\{L_{i}(0,d)=\ell\}}\mathbf{1}_{G^{i}_{0,d}}\mid\mathcal{F}_{i-1}]\leq 1.$
We now analyze the remaining case, i.e., when there are no defects in [iKK/2,iK+K/2)[iK-K/2,iK+K/2) after we stabilize $\eta^{\prime}_{i}(m,d)$ using the procedure described in Section 3. Thus, we need only to analyze the case where $d=0$.
We will use from now on a slight abuse of notation and denote by $L_{i}(j),S_{i}(j)$ and $\mathbf{1}_{G^{i}_{j}}$ the values of the quantities $L_{i},S_{i}$ and $\mathbf{1}_{G^{i}}$ after $j$ attempted emissions in block $i$.
For each particle added at $iK+a$, at least one attempted emission will be made. The only other way an attempted emission may be triggered is when a particle reaches the vacant hole at $iK$. In other words, for each $m$ there is at least one corresponding $j$. Therefore,
$\sum_{m=0}^{\infty}e^{48S_{i}(m,d)}\mathbf{1}_{\{L_{i}(m,d)=\ell\}}\mathbf{1}_{G^{i}_{m,d}}\leq\sum_{j=0}^{\infty}e^{48S_{i}(j)}\mathbf{1}_{\{L_{i}(j)=\ell\}}\mathbf{1}_{G^{i}_{j}}.$
@@
$\sup_{\ell\geq 0}\mathbb{E}\Bigg[\sum_{j=0}^{\infty}e^{48S_{i}(j)}\mathbf{1}_{\{L_{i}(j)=\ell\}}\mathbf{1}_{G^{i}_{j}}\mid\mathcal{F}_{i-1}\Bigg]\leq e^{4}.$ (2)
@@
It is immediate that $H(j)$ is $\mathcal{G}_{j}$-measurable. For brevity we will define $\tilde{\mathbb{P}}[\cdot]:=\mathbb{P}[\cdot\mid\mathcal{F}_{i-1}]$ and $\tilde{\mathbb{E}}[\cdot]:=\mathbb{E}[\cdot\mid\mathcal{F}_{i-1}]$. To prove Proposition 6 we will make use of the following lemmas.
Lemma 1.
@@
$\tilde{\mathbb{P}}[H(j)>a/2\mid\mathcal{G}_{j-1}]\mathbf{1}_{\{H(j-1)\in[0,a/2]\cup\{a\}\}}\leq e^{-150},$ (3)
$\tilde{\mathbb{P}}(a/2<H(j)<a\mid\mathcal{G}_{j-1})<e^{-150}.$ (4)
Remark 1.
@@
Lemma 1 says that, if $K$ is large enough, we can make the hole process to have a drift to the left.
Lemma 2.
@@
$\tilde{\mathbb{P}}(L(j+2)>L(j)\mid\mathcal{G}_{j})\geq\frac{1}{5}.$
For every integer $\ell\geq 0$, denote by $\tau_{\ell}$ the first $j$ for which $L(j)=\ell$, with $\tau_{-1}:=-1$. We have that $\tau_{\ell+1}>\tau_{\ell}$, because every attempted emission will result in at most one particle emitted to the left. We can rewrite (2) to obtain
$\mathbb{E}$ $\left[\sum_{j=\tau_{\ell-1}+1}^{\infty}e^{48S(j)}\mathbf{1}_{\{L(j)=\ell\}}\mathbf{1}_{G_{j}}\mid\mathcal{F}_{i-1}\right]$
$\leq\sum_{j=\tau_{\ell-1}+1}^{\infty}\tilde{\mathbb{E}}[\mathbf{1}_{\{L(j)=\ell\}}\mathbf{1}_{\{S(j)=0\}}+e^{48}\mathbf{1}_{\{L(j)=\ell\}}\mathbf{1}_{G_{j}}\mathbf{1}_{\{S(j)=1\}}]$
$\leq\tilde{\mathbb{E}}[\tau_{\ell+1}-\tau_{\ell}]+\sum_{k=1}^{\infty}\tilde{\mathbb{E}}[e^{48}\mathbf{1}_{\{L(\tau_{\ell-1}+k)=\ell\}}\mathbf{1}_{G_{\tau_{\ell-1}+k}}\mathbf{1}_{\{S(\tau_{\ell-1}+k)=1\}}].$ (5)
Using Lemma 2, we get
$\tilde{\mathbb{E}}[\tau_{\ell+1}-\tau_{\ell}]=\sum_{r=0}^{\infty}\tilde{\mathbb{P}}(\tau_{\ell+1}-\tau_{\ell}>r)\leq\sum_{r=0}^{\infty}\left(\frac{4}{5}\right)^{\lfloor r/2\rfloor}<10.$ (6)
We split the sum in (5) at $k_{0}:=250$. By Lemma 2,
$\sum_{k=k_{0}}^{\infty}$ $\tilde{\mathbb{E}}[e^{48}\mathbf{1}_{\{L(\tau_{\ell-1}+k)=\ell\}}\mathbf{1}_{G_{\tau_{\ell-1}+k}}\mathbf{1}_{\{S(\tau_{\ell-1}+k)=1\}}]$
$\leq\sum_{k=k_{0}}^{\infty}e^{48}\tilde{\mathbb{P}}(L(\tau_{\ell-1}+k)=\ell)\leq\sum_{k=k_{0}}^{+\infty}e^{48}\left(\frac{4}{5}\right)^{\lfloor k/2\rfloor}<1.$ (7)
@@
$\tilde{\mathbb{E}}[\mathbf{1}_{\{L(\tau_{\ell-1}+k)=\ell\}}\mathbf{1}_{G_{\tau_{\ell-1}+k}}\mathbf{1}_{\{S(\tau_{\ell-1}+k)=1\}}]<k_{0}^{-1}e^{-48}.$ (8)
Let us analyze first the case $\ell\geq 1$. In this case, obviously $\mathbf{1}_{G_{\tau_{\ell-1}}}=1$. First, we consider $k=1$. Note that if $S(\tau_{\ell-1}+1)=1$ and $L(\tau_{\ell-1}+1)=\ell$, then the $(\tau_{\ell-1}+1)$-th attempted emission is successful and $H(\tau_{\ell-1})=a$. Using (3), after conditioning on $\mathcal{G}_{\tau{\ell-1}}$,
$\tilde{\mathbb{E}}$ $[\mathbf{1}_{\{L(\tau_{\ell-1}+1)=\ell\}}\mathbf{1}_{\{S(\tau_{\ell-1}+1)=1\}}]$
$\leq\tilde{\mathbb{E}}[\mathbf{1}_{\{L(\tau_{\ell-1}+1)=\ell\}}\mathbf{1}_{\{H(\tau_{\ell-1}+1)=a\}}\mathbf{1}_{\{H(\tau_{\ell-1})=a\}}]$